# Tidal heating in binary inspiral of strange quark stars

### v1

@@ II.3 Bulk viscosity in strange quark stars
Tidal perturbations induce volume expansion and compression of the medium, which subject the system to instantaneous deviations from chemical equilibrium. For three-flavour quark matter in the normal phase where the medium is transparent to neutrinos, the non-leptonic weak process
@@
provide the main contribution to dissipate energy and re-establish chemical equilibrium at low temperatures $T\leq 10^{10}$a’d et al. (2007); Alford et al. (2010); Cruz Rojas et al. (2024); Hernández et al. (2024), this being the relevant regime for the tidal heating in strange quark stars.
@@
where $G_{F}=1.166\times 10^{-5}\,\text{GeV}^{-2}$ is the Fermi coupling constant, and $\Theta_{C}$ is the Cabibbo angle, and $k_{f}\equiv\sqrt{\mu_{f}^{2}-m_{f}^{2}}$.
Figure 3: Relative strengths of various sources of viscosities of dense matter inside neutron stars as a function of temperature at $n_{B}=2.5n_{0}$ with $n_{0}\approx 0.15$ fm-3. The bulk viscosity at $\omega=2\pi$ kHz before and after the resonant peak is shown for the strange-quark-matter EoSs (Set A-F) detailed in table 1, non-leptonic reactions involving $\Lambda$ hyperons, and m-Urca processes in nuclear matter. We also plot, the standard neutron matter shear viscosity coming from $ee$ scattering $\eta_{SV}^{ee}$, and the shear viscosity in degenerate quark matter coming from quark-quark scattering $\eta_{SV}^{qq}$.
The expression in Eq (20) considers part of the quark mass dependence of the rates. However, for high values of the strange quark mass the rate difference is a complex expression that can only be computed numerically Madsen (1993); Anand et al. (1997). Further, Eq. (20) expanded for $m_{s}=0$ and for massless light quarks at $\mu_{d}=\mu_{u}$ allow us to get the well-known result Madsen (1993)
@@
In Fig. 3, we display the bulk viscosity as a function of temperature at the typical characteristic value of oscillation frequency of $1$ kHz Pradhan and Chatterjee (2021) and $n_{B}=2.5n_{0}$, being $n_{0}$ the nuclear saturation density. First, we plot the bulk viscosity employing the EoSs described in table 1. Further, we include the bulk viscosity from non-leptonic reactions involving $\Lambda$-hyperons (Ghosh et al., 2024) as well as modified-Urca (m-Urca) reactions to compare their relative strengths. Lastly, we plot the shear viscosity from $ee$ scattering $\eta_{SV}^{ee}$ considering the Fermi-liquid approximation Flowers and Itoh (1976) and the shear viscosity in degenerate quark matter from quark-quark scattering $\eta_{SV}^{qq}$ at approximated equal quark chemical potentials $\mu_{q}=350$ MeV obtained at $2.5n_{0}$ and $\alpha_{s}\approx 0.7$ Heiselberg and Pethick (1993). At $T\sim 10^{9}$ K, bulk viscosity in unpaired strange quark matter reaches its resonant maximum value ($\sim 10^{29}$ gm cm-1s-1), which is several orders of magnitude higher than the $ee$ shear, m-Urca bulk viscosity Sawyer (1989b) and the $qq$ shear viscosity in this temperature regime. Regarding the bulk viscosity in strange quark matter, the values associated with the sets B, D and F (for which $m_{s}=200$ MeV) are a few times higher than the ones of sets A, C and E (with $m_{s}=150$ MeV). This trend is expected given the strong $m_{s}$-dependence of $\zeta$ in Eq. (22). On the other hand, the bulk viscosity does not depend on the bag constant since in this model there is neither temperature nor chemical potential dependence on $B_{\rm eff}$. Finally, its dependence on $a_{4}$ is almost negligible below $T=10^{9}$ K and small for higher temperatures up to $10^{11}$ K.
III Temperature estimate during binary inspiral
Since the dissipated tidal energy is converted to thermal energy and heats up the system during the inspiral, it is crucial to obtain an estimate of the average temperature we expect from this tidal dissipation. During the binary inspiral, the thermal evolution of the strange quark star can be written as Lai (1994)
@@ III Temperature estimate during binary inspiral
where $\dot{E}_{\rm cool}$ denotes the rate of cooling due to neutrino emission and surface photon luminosity. As shown in Appendix B, the cooling rate due to neutrino emission from strange quark stars is much larger than the inspiral timescale (of the order of $\sim 100$ s) for the relevant range of temperatures $\leq 10^{9}$ K. Therefore, the cooling due to neutrino emission can be neglected in this calculation. Then, the rate of change of the internal energy is given by
$\frac{dU}{dt}=C_{V}\frac{dT}{dt},$ (26)
@@
where $C_{V}$ is the specific heat capacity of the strange quark matter.
The heat capacity at constant volume and chemical potentials can be obtained from the thermodynamic potential in Eqs. (10) and (11) following a similar procedure as in Ref. Ferrer et al. (2021) assuming $\mu_{i}>0$ and $m_{i}<\mu_{i}$ with $i=u,d,s,e$ in the low temperature limit.
In table 3 we provide the ratio $\sigma=\bar{C}_{V,\text{ni}}/\bar{C}_{V,\text{F}}$ between the average heat capacity computed with the non-ideal bag model $\bar{C}_{V,\text{ni}}$, and the free Fermi gas approximation $\bar{C}_{V,\text{F}}$. We consider the average for a given density profile within the low temperature limit. Under these conditions, $\sigma$ slightly varies with the masses of the stars and temperature, then we only consider its variation with respect to the sets of parameters in table 1.
$\sigma$
Set A $0.9241$
Set B $0.9231$
Set C $0.8812$
Set D $0.8801$
Set E $0.8340$
Set F $0.8330$
Table 3: Ratio $\sigma=\bar{C}_{V,\text{ni}}/\bar{C}_{V,\text{F}}$ of the average heat capacities with the non-ideal bag model and the free Fermi gas using the sets in table 1.
@@
Given the temperature dependence of the bulk viscosity in unpaired strange quark matter described in Sec. II.3 (see also Eq. (20)), the bulk viscous dissipation rate can be parametrized as (see Eq. (7))
$\gamma=\frac{AT^{2}}{B+T^{4}}.$ (27)
The parameters $A$ and $B$ are numerically fitted to the functional dependence of $\gamma$ profiles with the temperature. After integrating the thermal evolution given in Eq. (25) from $D\to\infty$, when the stars were far apart and at a very low temperature $T_{0}\sim 10^{5}-10^{6}$ K, we can get an estimate of the temperature reached as a function of their separation $D$
$\begin{split}\frac{T^{4}}{4}+B\ln{\left(\frac{T}{T_{0}}\right)}&=\frac{T_{0}^{4}}{4}+\frac{\pi}{21870}\frac{\sigma Q^{2}q}{\bar{\omega}^{4}}\frac{A}{10^{22}}\frac{R^{3}}{M}\left(\frac{3R}{D}\right)^{5},\end{split}$ (28)
where the values of $\sigma$ are depicted in table 3 for the EoSs considered.
@@
Figure 4: Estimates of the average temperature due to the tidal dissipation as a function of GW frequency for the six different EoSs (Sets A-F) considered for three different equal mass binary systems: a) $1.4M_{\odot}$ (top panel) b) $1.6M_{\odot}$ (mid panel) and c) $1.8M_{\odot}$ (bottom panel).
@@
Fig. 4 shows the estimate of the average temperature of the star for the EoSs considered in this study and for three different equal mass binary neutron star systems. For all scenarios considered, we found temperatures of few times $10^{9}$ K reached at the end of the inspiral with a slight increase towards high mass binaries. In addition, as the strange quark mass increases from $150$ MeV (sets A, C and E) to $200$ MeV (sets B, D and F) the final temperature also increases around $25\%$ for the three stellar configurations. This is because the bulk viscosity increases with the strange quark mass. Moreover, we see a slight decrease in the temperature as we decrease $a_{4}$ and $B_{\rm eff}$ at fixed strange quark mass.
IV Phase Contribution to gravitational waveform
@@
During the early stages of a binary neutron star evolution away from the merger, the change in the orbital frequency due to the emission of GWs is much smaller than the orbital frequency itself. In this regime, the evolution of the orbital phase $\Phi(t)$ of the binary system is computed as a perturbative expansion in a small parameter, typically taken to be the characteristic velocity $v=(\pi Mf)^{1/3}$, $M$ being the total mass of the binary $M=M_{1}(1+q)$ and $f=\Omega/\pi$ being the GW frequency. This analytical procedure requires $v\ll 1$, which makes it useful in the early inspiral phase of the binary. In this adiabatic regime, the loss of binding energy $E(v)$ of the two-body system with time equals the GW flux emitted to future null infinity $\mathcal{F}^{\infty}(v)$ plus the energy dissipated to thermal energy due to the viscous dissipation $\dot{E}_{\rm visc}(v)$. Then, the energy balance condition becomes
@@
$\dot{E}_{\rm visc}=\frac{24\pi}{5}q^{2}(1+q)\frac{M_{1}^{2}}{R}\frac{Q^{2}}{\bar{\omega}^{4}}\left(\frac{R}{M}\right)^{9}\gamma(v)\,v^{18}\,.$ (35)
Note that we express $\gamma$ as a function of the velocity simply because the magnitude of the viscosity strength depends strongly on the temperature, and due to viscous dissipation we expect the star also to heat up during inspiral. For equal mass binary systems considered in this work, we set $q=1$ and also multiply the expression in Eq. (35) by a factor of $2$ to consider the contribution of the both the binary components.
@@
where $\eta=qM_{1}^{2}/M^{2}$ is the symmetric mass ratio. For equal mass systems $\eta=1/4$. Plugging the expressions for $\dot{E}_{\rm visc}$, $E(v)$ and $\mathcal{F}^{\infty}(v)$ from Eqs. (35), (36) and (37) respectively in Eq. (32), we numerically integrate to obtain the phase of gravitational waves. To compute the additional phase due to viscous dissipation $\Delta\Phi$, we simply subtract this phase from the same obtained without considering the contribution from viscous dissipation $\dot{E}_{\rm visc}$.
In Fig. 5 we show this additional phase for the EoSs considered in this study and three different BNS systems of equal mass. For all these systems considered, the additional phase obtained is of the order of $\sim 0.1-0.5$ rad for the sets with $m_{s}=200$ MeV (sets B, D and F) and $0.05-0.1$ rad for the sets with $m_{s}=150$ MeV (sets A, C and E). As seen in the case of the temperature estimates as well as in Sec. II, this is because the rate of dissipated energy is proportional to the bulk viscous dissipation rate and therefore to the bulk viscosity. Thus, we expect that $\dot{E}_{\rm visc}$ increases as the strange quark mass increases. We also see a slight increase in the phase difference towards high mass binaries for all the EoSs and also for decreasing $a_{4}$ and bag constant $B_{\rm eff}^{1/4}$ for all the equal mass binaries.
As discussed in  (Read, 2023), for this phase shift in the gravitational waveform to be detectable by current or next generation ground-based GW detectors, it should be smaller than the tolerance limit
$|\Delta\Phi(f)|\leq\frac{\sqrt{S_{n}(f)}}{2A(f)\sqrt{f}}\leavevmode\nobreak\ ,$ (38)
@@
where $S_{n}(f)$ is the power spectral density of the noise for each detector and $A(f)$ is the gravitational wave amplitude of the waveform. Current waveform uncertainties at frequencies much lower than the merger frequency are much lower than the estimated phase uncertainties we are interested in here. So, we only consider the amplitude from a single state-of-the-art waveform model ‘IMRPhenomPv2_NRTidalv2’ for this analysis  (Ajith et al., 2007; Husa et al., 2016; Khan et al., 2016; Hannam et al., 2014; Dietrich et al., 2017, 2019). In Fig. 5, we plot this estimated phase sensitivity according to Eq. (38) for the detector sensitivity of LIGO A+ (LVK, 2022), 40-km Cosmic Explorer (CE) (Evans et al., 2021; Reitze et al., 2019) and the Einstein Telescope (ET)-D (et all., 2011), at a distance of 100 Mpc. From these estimates, we see that this extra phase shift should in principle be detectable with next-generation detectors for nearby sources with high enough signal-to-noise (SNR) ratio.
V Discussion

### v2

@@ III.3 Bulk viscosity in strange quark stars
Tidal perturbations induce volume expansion and compression of the medium, which subject the system to instantaneous deviations from chemical equilibrium. For three-flavour quark matter in the normal phase, where the medium is transparent to neutrinos, the non-leptonic weak process
@@
provides the main contribution to dissipate energy and re-establish chemical equilibrium at low temperatures ($T\leq 10^{10}$ K) Sa’d et al. (2007); Alford et al. (2010); Cruz Rojas et al. (2024); Hernández et al. (2024). This temperature regime is relevant for the tidal heating in strange quark stars.
@@
where $G_{F}=1.166\times 10^{-5}\,\text{GeV}^{-2}$ is the Fermi coupling constant, $\Theta_{C}$ is the Cabibbo angle, and $k_{f}\equiv\sqrt{\mu_{f}^{2}-m_{f}^{2}}$. The expression in Eq (21) considers part of the quark mass dependence of the rates. However, for high values of the strange quark mass the rate difference is a complex expression that can only be computed numerically Madsen (1993); Anand et al. (1997). Further, Eq. (21) expanded for $m_{s}=0$ and for massless light quarks at $\mu_{d}=\mu_{u}$ allow us to get the well-known result Madsen (1993)
@@
Figure 3: Relative strengths of various sources of viscosities of dense matter inside neutron stars as a function of temperature at $n_{B}=2.5n_{0}$ with $n_{0}\approx 0.15$ fm-3. The bulk viscosity at $\omega=2\pi$ kHz before and after the resonant peak is shown for the strange-quark-matter EoSs (Set A-F) detailed in table 1, non-leptonic reactions involving $\Lambda$ hyperons, and m-Urca processes in nuclear matter. We also plot, the standard neutron matter shear viscosity coming from $ee$ scattering $\eta_{SV}^{e}$, and the shear viscosity in degenerate quark matter coming from quark-quark scattering $\eta_{SV}^{qq}$.
In Fig. 3, we display the bulk viscosity as a function of temperature at the typical characteristic value of oscillation frequency of $1$ kHz Pradhan and Chatterjee (2021) and $n_{B}=2.5n_{0}$, being $n_{0}$ the nuclear saturation density. First, we plot the bulk viscosity employing the EoSs described in table 1. Further, we include the bulk viscosity from non-leptonic reactions involving $\Lambda$-hyperons (Ghosh et al., 2024) as well as modified-Urca (m-Urca) reactions to compare their relative strengths. Lastly, we plot the neutron matter shear viscosity from electron scattering $\eta_{SV}^{e}$ according to Alford et al. (2012) at a proton fraction $x_{p}=0.1$ and $3n_{0}$ and the shear viscosity in degenerate quark matter from quark-quark scattering $\eta_{SV}^{qq}$ at approximated equal quark chemical potentials $\mu_{q}=350$ MeV obtained at $2.5n_{0}$ and $\alpha_{s}\approx 0.7$ Heiselberg and Pethick (1993). At $T\sim 10^{9}$ K, bulk viscosity in unpaired strange quark matter reaches its resonant maximum value ($\sim 10^{29}$ g cm-1s-1), which is several orders of magnitude higher than the $ee$ shear, m-Urca bulk viscosity Sawyer (1989b) and the $qq$ shear viscosity in this temperature regime. Regarding the bulk viscosity in strange quark matter, the values associated with the sets B, D and F (for which $m_{s}=200$ MeV) are a few times higher than the ones of sets A, C and E (with $m_{s}=150$ MeV). This trend is expected given the strong $m_{s}$-dependence of $\zeta$ in Eq. (23). On the other hand, the bulk viscosity does not depend on the bag constant since in this model there is neither temperature nor chemical potential dependence on $B_{\rm eff}$. Finally, its dependence on $a_{4}$ is almost negligible below $T=10^{9}$ K and small for higher temperatures up to $10^{11}$ K.
IV Temperature estimate during binary inspiral
Since the dissipated tidal energy is converted to thermal energy and heats up the system during the inspiral, it is crucial to obtain an estimate of the average temperature we expect from this tidal dissipation. In this study, we consider equal mass binaries, and hence we remove the subscript associated with each star from here on. Considering the high thermal conductivity, the stars are expected to be isothermal during the inspiral. Considering the effects of General Relativity, the constant redshifted temperature, $\tilde{T}$ relates to the local temperature, $T$ by
$\tilde{T}=Te^{\nu},$ (26)
where $\nu$ is the metric function (see Eq. (48)), which determines the gravitational redshift. During the binary inspiral, the thermal evolution for the redshifted temperature of the strange quark star can be written as Lai (1994)
@@ IV Temperature estimate during binary inspiral
where $U$ is the internal energy of the system, $\dot{E}_{\rm cool}$ denotes the rate of cooling due to neutrino emission and surface photon luminosity. As shown in Appendix B, the cooling timescale due to neutrino emission from strange quark stars is much longer than the inspiral timescale (of the order of $\sim 100$ s) for the relevant range of temperatures $\leq 10^{9}$ K. Therefore, the cooling due to neutrino emission can be neglected in this calculation. In addition, photon emission in quark stars has been studied in Ref. Cheng and Harko (2003). This results to be smaller than the blackbody radiation and not relevant for the temperatures we consider in this work. Thus, the rate of change of the internal energy is given by
$\frac{dU}{dt}=C_{V}\frac{d\tilde{T}}{dt},$ (28)
where $C_{V}$ is the average heat capacity of strange quark matter. This quantity can be computed over the entire stellar volume following a similar reasoning as in Ref. Ofengeim et al. (2017)
$C_{V}=\int_{0}^{R}c_{V}(\rho,\tilde{T})\frac{4\pi r^{2}dr}{\sqrt{1-2m/r}}.$ (29)
Here $R$ is the radius of the star as stated above, and $m=m(r)$ is the gravitational mass inside a sphere with circumferential radius $r$.
@@
The specific heat capacity, $c_{V}$ can be computed in strange quark matter using the following thermodynamic relation Ferrer et al. (2021)
$c_{V}=-T\frac{\partial^{2}\Omega_{\rm ni}}{\partial T^{2}}\Bigg|_{\mu_{i}},$ (30)
with $i=u,d,s,e$. Hereafter, we use the low-temperature approximation $T<\mu_{i}$ and $m_{i}<\mu_{f}$. These approximations are valid for the chemical potentials and temperatures considered for electric charge neutral strange quark matter in chemical equilibrium. Following a similar procedure as the one to obtain Eq. (11) in Ref. Ferrer et al. (2021), we get the specific heat capacity of strange quark matter
$c_{V}=\tilde{T}e^{-\nu}\left(\frac{\mu_{e}^{2}}{3}+a_{4}\sum_{f=u,d,s}\mu_{f}\sqrt{\mu_{f}^{2}-m_{f}^{2}}\right).$ (31)
Note that with the definitions above and considering a constant redshifted temperature over the stellar volume, we can express the average heat capacity as follows
$C_{V}=\sigma\tilde{T}$, (32)
where $\sigma$ depends on the variation of the chemical potentials, the gravitational redshift $\nu$ profile, and also the mass configurations.
In table 3 we provide the values of $\sigma$ using the non-ideal bag model for the different equal-mass configurations considered in this work.
Set $\sigma$ $\left[\text{erg}/\text{K}^{2}\right]$ $\sigma$ $\left[\text{erg}/\text{K}^{2}\right]$ $\sigma$ $\left[\text{erg}/\text{K}^{2}\right]$
$(M/M_{\odot}=1.4)$ $(M/M_{\odot}=1.6)$ $(M/M_{\odot}=1.8)$
A $9.641\times 10^{29}$ $1.125\times 10^{30}$ $1.295\times 10^{30}$
B $9.578\times 10^{29}$ $1.118\times 10^{30}$ $1.288\times 10^{30}$
C $9.615\times 10^{29}$ $1.119\times 10^{30}$ $1.283\times 10^{30}$
D $9.544\times 10^{29}$ $1.112\times 10^{30}$ $1.275\times 10^{30}$
E $9.517\times 10^{29}$ $1.108\times 10^{30}$ $1.267\times 10^{30}$
F $9.466\times 10^{29}$ $1.099\times 10^{30}$ $1.257\times 10^{30}$
Table 3: $\sigma$ values in the average heat capacity ($C_{V}=\sigma\tilde{T}$) for different masses and EoSs considered in this work.
@@
Given the local-temperature dependence of the bulk viscosity in unpaired strange quark matter described in Sec. III.3 (see also Eq. (21)), the bulk viscous dissipation rate can be parametrized as (see Eq. (8))
$\gamma=\frac{A\tilde{T}^{2}}{B+\tilde{T}^{4}}.$ (33)
The parameters $A$ and $B$ are numerically fitted to the $\gamma$ profiles with the isothermal temperature corrected for the gravitational redshift.
@@
Figure 4: Estimates of the average final temperature due to the tidal dissipation as a function of GW frequency for the six different EoSs (Sets A-F) considered for three different equal mass binary systems: a) $1.4M_{\odot}$ (top panel) b) $1.6M_{\odot}$ (mid panel) and c) $1.8M_{\odot}$ (bottom panel).
Then, we obtain an estimate of the temperature reached, $\tilde{T}_{f}$, as a function of the separation of the stars $D$. Using Eq. (10) with the orbital evolution equation
$\frac{dD}{dt}=-\frac{64}{5}\frac{M_{1}^{3}q(1+q)}{D^{3}},$ (34)
we integrate the thermal evolution equation in Eq. (27) from $D\to\infty$, when the stars are far apart and at a very low temperature $\tilde{T}_{0}\sim 10^{5}-10^{6}$ K, to finally obtain
[…]
@@
Fig. 4 shows the estimate of the average temperature, $\tilde{T}_{f}$, of the star for the EoSs considered in this study and for three different equal mass binary neutron star systems. For all scenarios considered, we found temperatures of a few times $10^{9}$ K reached at the end of the inspiral with a slight increase towards high mass binaries. In addition, as the strange quark mass increases from $150$ MeV (sets A, C and E) to $200$ MeV (sets B, D and F) the final temperature also increases around $33\%$ for the three stellar configurations. This is because the bulk viscosity increases with the strange quark mass. Moreover, we see a slight decrease in the temperature as we decrease $a_{4}$ and $B_{\rm eff}$ at fixed strange quark mass.
V Phase Contribution to gravitational waveform
@@
During the early stages of a binary neutron star evolution away from the merger, the change in the orbital frequency due to the emission of GWs is much smaller than the orbital frequency itself. In this regime, the evolution of the orbital phase $\Phi(t)$ of the binary system is computed as a perturbative expansion in a small parameter, typically taken to be the characteristic velocity $v=(\pi Mf)^{1/3}$,$M$ being the total mass of the binary and $f=\Omega/\pi$ being the GW frequency. This analytical procedure requires $v\ll 1$, which makes it useful in the early inspiral phase of the binary. In this adiabatic regime, the loss of binding energy $E(v)$ of the two-body system with time equals the GW flux emitted to future null infinity $\mathcal{F}^{\infty}(v)$ plus the energy dissipated to thermal energy due to the viscous dissipation $\dot{E}_{\rm visc}(v)$. Then, the energy balance condition becomes
@@
$\dot{E}_{\rm visc}=\frac{24\pi}{5}q^{2}(1+q)\frac{(M/2)^{2}}{R}\frac{Q^{2}}{\bar{\omega}^{4}}\left(\frac{R}{M}\right)^{9}\gamma(v)\,v^{18}\,.$ (42)
Note that we express $\gamma$ as a function of the velocity simply because the magnitude of the viscosity strength depends strongly on the temperature, and due to viscous dissipation we expect the star also to heat up during inspiral. Also, for equal mass binary systems considered in this work, we set $q=1$ and also multiply the expression in Eq. (42) by a factor of $2$ to consider the contribution of the both the binary components.
@@
where $\eta=M_{1}M_{2}/M^{2}$ is the symmetric mass ratio. For equal mass systems $\eta=1/4$. Plugging the expressions for $\dot{E}_{\rm visc}$, $E(v)$ and $\mathcal{F}^{\infty}(v)$ from Eqs. (42), (43) and (44) respectively in Eq. (39), we numerically integrate it to obtain the phase of gravitational waves. To compute the additional phase due to viscous dissipation $\Delta\Phi$, we simply subtract this phase from the same obtained without considering the contribution from viscous dissipation $\dot{E}_{\rm visc}$. If we ignore the additional frequency dependence of the dissipation rate $\gamma$, analytic computation of Eq. (39) tells us that tidal dissipation effects enter the GW phase at 4th Post Newtonian (PN) order compared to the point-particle phase with an additional $\log(v)$ term making it non-degenerate with any other physical binary parameters, particularly time of coalescence Ripley et al. (2023).
@@
In Fig. 5 we show this additional phase for the EoSs considered in this study and three different BNS systems of equal mass. For all these systems considered, the additional phase obtained is of the order of $\sim 0.1-0.5$ rad for the sets with $m_{s}=200$ MeV (sets B, D and F) and 0.060.20.06-0.2 rad for the sets with $m_{s}=150$ MeV (sets A, C and E). As seen in the case of the temperature estimates as well as in Sec. II, this is because the rate of dissipated energy is proportional to the bulk viscous dissipation rate and therefore to the bulk viscosity. Thus, we expect that $\dot{E}_{\rm visc}$ increases as the strange quark mass increases. We also see a slight increase in the phase difference towards low mass binaries for all the EoSs and also for decreasing $a_{4}$ and bag constant $B_{\rm eff}^{1/4}$ for all the equal mass binaries.
As discussed in (Read, 2023), for this phase shift in the gravitational waveform to be detectable by current or next generation ground-based GW detectors, it should be larger than the tolerance limit
$|\Delta\Phi(f)|\geq\frac{\sqrt{S_{n}(f)}}{2\tilde{A}(f)\sqrt{f}}\penalty 10000\ ,$ (45)
where $S_{n}(f)$ is the power spectral density of the noise for each detector and $\tilde{A}(f)$ is the gravitational wave amplitude of the waveform. Current waveform uncertainties at frequencies much lower than the merger frequency are much lower than the estimated phase uncertainties we are interested in here. So, we only consider the amplitude from a single state-of-the-art waveform model ‘IMRPhenomPv2_NRTidalv2’ for this analysis (Ajith et al., 2007; Husa et al., 2016; Khan et al., 2016; Hannam et al., 2014; Dietrich et al., 2017, 2019). In Fig. 5, we plot this estimated phase sensitivity according to Eq. (45) for the detector sensitivity of LIGO A+ (LVK, 2022), 40-km Cosmic Explorer (CE) (Evans et al., 2021; Reitze et al., 2019) and the Einstein Telescope (ET)-D (et all., 2011), at a distance of 100 Mpc. From these estimates, we see that this extra phase shift should in principle be detectable with next-generation detectors for nearby sources with high enough signal-to-noise (SNR) ratio.
VI Discussion